OpenStudy (anonymous):

A 1200W kettle is filled with 0.5 kg of water at 28 degree Celcius, it is turned on until the water boils. What is the time taken to boil the water?

OpenStudy (anonymous):

From the equation of thermidometry we have Q=m*c*dT where m the mass of the water, c the heat capacity of water and dT the difference in temperature.We know that water boils in 100 celsius degrees so we have dT=72.So Q=0.5*4185.5*72=150678J.We know that the power of the kettle is 1200W.We have P=W/t => t=W/T <=> t=150678/1200, so we get t is approximately equal to 125.5s

OpenStudy (anonymous):

I forgot to mention that Q is the amount of heat the kettle supplies to the water.

OpenStudy (anonymous):

Pt = mctheta???

OpenStudy (anonymous):

yes, that's a rapid way to solve th problem.

OpenStudy (anonymous):

h=pt w=j but 1200=1200xt t=1200/1200 t=1s

OpenStudy (anonymous):

@pastory I think what you wrote doesn't make sense.

OpenStudy (anonymous):

How ever, i still havent get the answer Could you show me the working please??

OpenStudy (anonymous):

@maths911 the basic equation of thermidometry states that the amount of heat transfered from a heating source to an other body is equal to the product of the mass of the body,the capacity of the body(a specific quantity which can be experimentaly found and depends on the material) and the difference in temperature between the boiling point and the temperature of the body.Now from all these we get the thermal energy (aka heat) the source transfers to the body.Since we know the electric power of the source (aka kettle) we solve through t to find time needed to boil water.